i need help on 47,49,61
45-74. Definite integrals Use a change of variables or Table 2.0 TO evaluate the following definite integrals.
47. [ 2x(4 – x²) dx 13 2x 49. ), 2x + 4 dx
84 6 61. st x Jo x² + x - dx ax
Answer #1

GD Evaluate ( 22 (4x2) de Solution Testing the substitution 0-28 dn. da Differentiating with respect to the du = d (4-2) = 0.Ji 2944 VI 2914 ③ Evaluate (3 21 de solution: Let I= . 13. 24 dn taking the substitution [u= 24 Differentialing with respectlogica du z [ Logouston Felgens Mag (12) - log69] hanging l bygg )] - logom pulang . = logem 1902) 34 (1 61 Evaluate la o teer uanzel limit will be change accordingly at nzo, u= 0+1= 1 af =4, ua 43+1 = 16+1=17 ie the linist of integration is [u= 4.

If you have any doubt in the solution kindly mention in the comment section. I will assist you as soon as possible.